P8-20: Interpreting Beta. A firm wishes to assess the impact of changes in the market return on an asset that has a beta of 1.20.

A. If the market return increased by 15%, what impact would this change be expected to have on the asset's return?

B. If the market return is decreased by 8%, what impact would this change be expected to have on the asset's return?

C. If the market return did not change, what impact, if any, would be expected on the asset's return?

D. Would this asset be considered more or less risky than the market? explain.
Answer #1

A)Asset Return will increase by : 15 *1.20 = 18%

B)Asset return will decrease by : 8 *1.20 = 9.6%

c)Asset return will not change if market return did not change

d)The market beta is 1 .Since the beta of asset (1.2) is more than beta of market ,asset is more risky than market. It means with 1% change in market asset will change by 1*1.2 =1.2 %